21. Ethical behavior is a. always doing what benefits yourself regardless of the consequences to others affected by your deci25. Assume that the accounts receivable balances are $110,000, $112,000 and $116,000 for 2021, 2022, and 2023, respectively.
Answer #1

Answers:

21.  Answer : Option c - knowing right from wrong and conducting yourself accordingly so that your decisions are consistent with your own value system and the values of those affected by your decisions.

Explanation: As stated in Managerial Accounting by Charles E. Davis - Ethics is knowing right from wrong and conducting yourself accordingly so that your decisions are consistent with your own value system and the values of those affected by your decisions. According to Institute of Business Ethics - There are three tests on ethical business decisions: First one is 'Do i mind others knowing what i've done?', second one is 'Who does my decision affect?', and the third one is 'Would my decision be taken fair by those who have got affected by my decision?'. So in the option (a) and (b), it is not concerned whether the decision is affecting others badly, hence it is wrong and in option (d) it states that aim is to to benefit others affected by your decisions rather than yourself, which is wrong too as the aim is to be consistent with your own value systems and the values of others affected by your decision.

22. Answer : Option c - Horizontal analysis

Explanation: Horizontal analysis compares account balances over different time periods. Horizontal analysis involves the computation of amount changes and percentage changes reported on the financial statement. Option (a) and (b) is incorrect as Vertical analysis also called as Common size analysis shows each amount on financial statement as a percentage of another amount. Option (d) is incorrect as Benchmarking means comparing one firms performance with other businesses within their trade or an industry standard.

23. Answer: Option a -  looking at the changes in the account balances reported on a financial statement over time.

Explanation: As stated in the Question 22, Horizontal analysis compares account balances over different time periods which are reported in the financial statements.

24. Answer: Option d - 20%

Calculation : (120,000-100,000 = 20,000; = 20000/100,000 = 0.02*100 = 20%) Accounts recievable balance increased from beginning, hence answer is Option (d) 20% increase.

25. Answer: Option d - 105.5%

Calculation : $116,000 ÷ $110,000 = 105.5%